Expalin how you know that 7/12 is greater than 1/3 but less than 2/3?
Answer by TimothyLamb(4379)

convert all of these to a common denominator:
7/12 , 1/3 , 2/3
---
7/12 = 7/12
1/3 = 4/12
2/3 = 8/12
---
place into compound inequality:
---
(1/3=4/12) < (7/12=7/12) < (2/3=8/12)
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php